# MultiSocial: Multilingual Benchmark of Machine-Generated Text Detection of Social-Media Texts

### v1

@@ B.3 Post-processing of Generated Texts
Both, the human and machine texts are cleaned by removing leading and trailing white-spaces, removing characters making problems in Polyglot161616based on https://github.com/aboSamoor/polyglot/issues/71#issuecomment-707997790, truncating the parts of texts above 200 words, and dropping duplicates and the samples with less than 3 words. Thus, we ensure that no text sample has multiple labels and that both human and machine samples are processed in the same way.
@@ Appendix D Cross-domain Evaluation
In Table 12 and 13, the results are provided in the same way as in Table 3, while testing on MULTITuDE (news domain) data. In pure cross-domain evaluation (Table 12), the machine texts are generated by the same generators (as used for training), in out-of-distribution evaluation (Table 13), the machine texts are generated only by Llama-2-70b (not available in MultiSocial for training).
@@ Appendix F Results Data
Tables 16-18 contain per-generator AUC ROC performance of each MGT detection method for each test language, Tables 19-21 contain per-platform AUC ROC performance of each MGT detection method for each test language, and Tables 22-24 contains per-generator AUC ROC performance of each MGT detection method for each platform, separately for each MGTD category. Table 25 contains cross-lingual evaluation of differently fine-tuned MGT detectors. Table 26 contains cross-platform evaluation of differently fine-tuned MGT detectors.
@@
BLOOMZ-3b-mixed-Detector Aya-101 0.83 0.88 0.82 0.88 0.85 0.82 0.85 0.84 0.91 0.79 0.68 0.86 0.92 0.84 0.84 0.85 0.78 0.78 0.83 0.76 0.81 0.86 0.83
@@
ChatGPT-Detector-RoBERTa-Chinese Aya-101 0.75 0.79 0.73 0.63 0.75 0.70 0.82 0.75 0.78 0.61 0.52 0.61 0.72 0.67 0.61 0.61 0.67 0.73 0.60 0.60 0.71 0.76 0.67
@@
RoBERTa-large-OpenAI-Detector Aya-101 0.76 0.46 0.46 0.19 0.40 0.72 0.73 0.43 0.32 0.37 0.46 0.31 0.20 0.39 0.31 0.35 0.36 0.57 0.35 0.45 0.43 0.69 0.43
@@
ruRoBERTa-ruatd-binary Aya-101 0.50 0.61 0.57 0.47 0.48 0.45 0.60 0.54 0.48 0.47 0.49 0.53 0.56 0.57 0.48 0.50 0.43 0.64 0.52 0.57 0.50 0.45 0.53
@@
XLM-RoBERTa-large-MultiSocial Aya-101 0.89 0.97 0.94 0.97 0.95 0.94 0.93 0.94 0.95 0.81 0.73 0.94 0.98 0.95 0.95 0.91 0.93 0.90 0.96 0.87 0.86 0.85 0.93
@@
mDeBERTa-v3-base-MultiSocial Aya-101 0.88 0.97 0.92 0.97 0.95 0.93 0.94 0.94 0.97 0.86 0.74 0.96 0.99 0.95 0.95 0.93 0.93 0.89 0.95 0.89 0.87 0.88 0.93
@@
Fast-Detect-GPT Discord N/A N/A 0.64 0.81 0.81 N/A 0.86 0.82 0.70 0.76 0.76 0.81 0.79 0.81 0.83 0.84 0.80 N/A 0.65 N/A N/A N/A 0.79
@@
BLOOMZ-3b-mixed-Detector Discord N/A N/A 0.96 0.89 0.87 N/A 0.87 0.90 0.87 0.80 0.66 0.84 0.86 0.91 0.91 0.89 0.78 N/A 0.71 N/A N/A N/A 0.87
@@
ChatGPT-Detector-RoBERTa-Chinese Discord N/A N/A 0.88 0.72 0.86 N/A 0.95 0.88 0.85 0.73 0.62 0.66 0.82 0.71 0.70 0.75 0.83 N/A 0.68 N/A N/A N/A 0.77
@@
Longformer Detector Discord N/A N/A 0.25 0.41 0.47 N/A 0.68 0.44 0.40 0.50 0.46 0.50 0.40 0.42 0.44 0.36 0.33 N/A 0.36 N/A N/A N/A 0.44
@@
RoBERTa-large-OpenAI-Detector Discord N/A N/A 0.25 0.05 0.12 N/A 0.44 0.10 0.18 0.32 0.52 0.12 0.09 0.09 0.07 0.09 0.09 N/A 0.22 N/A N/A N/A 0.16
@@
ruRoBERTa-ruatd-binary Discord N/A N/A 0.51 0.41 0.25 N/A 0.53 0.42 0.35 0.41 0.36 0.45 0.43 0.43 0.40 0.35 0.25 N/A 0.32 N/A N/A N/A 0.40
@@
BLOOMZ-3b-MultiSocial Discord N/A N/A 0.99 1.00 0.97 N/A 0.99 0.99 0.99 0.97 0.84 0.98 1.00 0.98 0.99 0.99 0.99 N/A 0.97 N/A N/A N/A 0.99
@@
Falcon-rw-1b-MultiSocial Discord N/A N/A 0.99 1.00 0.97 N/A 0.99 0.98 0.99 0.96 0.87 0.98 1.00 0.99 0.99 0.99 0.99 N/A 0.97 N/A N/A N/A 0.99
@@
Llama-3-8b-MultiSocial Discord N/A N/A 0.99 1.00 0.97 N/A 0.99 0.99 1.00 0.96 0.88 0.99 1.00 0.99 1.00 0.99 0.99 N/A 0.98 N/A N/A N/A 0.99
@@
Mistral-7b-v0.1-MultiSocial Discord N/A N/A 0.99 1.00 0.97 N/A 0.99 0.99 1.00 0.96 0.91 0.99 1.00 0.99 1.00 0.99 0.99 N/A 0.98 N/A N/A N/A 0.99
@@
XLM-RoBERTa-large-MultiSocial Discord N/A N/A 0.96 0.99 0.96 N/A 0.98 0.97 0.99 0.89 0.72 0.98 1.00 0.98 0.99 0.97 0.98 N/A 0.99 N/A N/A N/A 0.97
@@
mDeBERTa-v3-base-MultiSocial Discord N/A N/A 0.98 1.00 0.97 N/A 0.98 0.98 0.99 0.91 0.75 0.98 1.00 0.98 0.99 0.99 0.99 N/A 0.97 N/A N/A N/A 0.98
@@
Binoculars Mistral-7B-Instruct-v0.2 0.77 0.62 0.68 0.68 0.67 0.68
@@
DetectLLM-LRR Mistral-7B-Instruct-v0.2 0.94 0.63 0.71 0.71 0.68 0.71
@@
Fast-Detect-GPT Mistral-7B-Instruct-v0.2 0.65 0.51 0.58 0.59 0.62 0.58
@@
LLM-Deviation Mistral-7B-Instruct-v0.2 0.95 0.63 0.69 0.70 0.67 0.70
@@
BLOOMZ-3b-mixed-Detector Mistral-7B-Instruct-v0.2 0.87 0.62 0.76 0.70 0.78 0.74
@@
ChatGPT-Detector-RoBERTa-Chinese Mistral-7B-Instruct-v0.2 0.85 0.70 0.75 0.74 0.85 0.76
@@
Longformer Detector Mistral-7B-Instruct-v0.2 0.51 0.55 0.59 0.44 0.47 0.52
@@
RoBERTa-large-OpenAI-Detector Mistral-7B-Instruct-v0.2 0.11 0.38 0.30 0.38 0.33 0.31
@@
ruRoBERTa-ruatd-binary Mistral-7B-Instruct-v0.2 0.47 0.52 0.54 0.53 0.53 0.53
@@
Aya-101-MultiSocial Mistral-7B-Instruct-v0.2 1.00 0.98 0.99 0.99 1.00 0.99
@@
BLOOMZ-3b-MultiSocial Mistral-7B-Instruct-v0.2 1.00 0.94 0.97 0.98 0.99 0.97
@@
Falcon-rw-1b-MultiSocial Mistral-7B-Instruct-v0.2 1.00 0.93 0.97 0.98 0.98 0.97
@@
Llama-3-8b-MultiSocial Mistral-7B-Instruct-v0.2 1.00 0.98 0.99 0.99 1.00 0.99
@@
Mistral-7b-v0.1-MultiSocial Mistral-7B-Instruct-v0.2 1.00 0.97 0.99 0.99 1.00 0.99
@@
XLM-RoBERTa-large-MultiSocial Mistral-7B-Instruct-v0.2 0.99 0.95 0.97 0.97 0.99 0.97
@@
mDeBERTa-v3-base-MultiSocial Mistral-7B-Instruct-v0.2 0.99 0.94 0.98 0.98 0.99 0.97
@@
XLM-RoBERTa-large Discord 0.99 0.87 0.90 0.83 0.94 0.90
@@
mDeBERTa-v3-base Discord 0.99 0.86 0.89 0.84 0.93 0.89

### v2

@@ B.3 Post-processing of Generated Texts
Both, the human and machine texts are cleaned by removing leading and trailing white-spaces, removing characters making problems in Polyglot161616based on https://github.com/aboSamoor/polyglot/issues/71#issuecomment-707997790, truncating the parts of texts above 200 words, and dropping duplicates and the samples with less than 3 words. Thus, we ensure that no text sample has multiple labels and that both human and machine samples are processed in the same way (avoiding processing bias of the detection).
@@
B.4 Meta-evaluation of Text Quality
Metric Meta-evaluator AR BN EN FR HI JA RU SW TR ZH $\rightarrow$ Average
Linguistic
Acceptability
Meta-Llama-3.1-70B-Instruct 0.74 0.07 0.65 0.68 0.68 0.53 0.57 0.55 0.66 0.87 0.60
Phi-3.5-mini-instruct 0.75 0.21 0.64 0.71 0.70 0.55 0.41 0.43 0.61 0.81 0.58
Qwen2-72B-Instruct 0.74 0.20 0.65 0.81 0.58 0.43 0.74 0.46 0.63 0.87 0.61
Aya-23-35B 0.72 0.19 0.59 0.80 0.66 0.51 0.43 0.38 0.59 0.86 0.57
Gemma-2-27b-it 0.72 0.25 0.60 0.68 0.68 0.56 0.70 0.80 0.67 0.85 0.65
GPT-4 0.71 0.22 0.82 0.81 0.61 0.47 0.80 0.76 0.72 0.85 0.68
Output Content Quality
Meta-Llama-3.1-70B-Instruct 0.70 0.05 0.64 0.71 0.71 0.54 0.77 0.64 0.58 0.85 0.62
Phi-3.5-mini-instruct 0.70 0.23 0.65 0.73 0.63 0.52 0.48 0.44 0.30 0.87 0.56
Qwen2-72B-Instruct 0.70 0.34 0.63 0.73 0.69 0.55 0.84 0.66 0.55 0.87 0.66
Aya-23-35B 0.66 0.15 0.57 0.68 0.65 0.56 0.43 0.37 0.27 0.89 0.52
Gemma-2-27b-it 0.71 0.35 0.62 0.69 0.63 0.49 0.89 0.84 0.69 0.77 0.67
GPT-4 0.69 0.26 0.68 0.72 0.65 0.51 0.92 0.88 0.68 0.84 0.68
Table 13: Correlation of open LLMs meta-evaluation of text quality with human judgements using METAL dataset. The reported values represent pairwise agreement using weighted F1-score, analogously to the detailed prompting strategy in Table 3 of (Hada et al., 2024). GPT-4 values are taken from the METAL dataset.
Linguistic Acceptability $\uparrow$ Output Content Quality $\uparrow$
Approach Meta-Llama-3.1-70B-Instruct Qwen2-72B-Instruct Gemma-2-27b-it Meta-Llama-3.1-70B-Instruct Qwen2-72B-Instruct Gemma-2-27b-it $\rightarrow$ Average
k_to_one 0.41 0.79 0.24 0.28 0.21 0.10 0.34
keywords 0.40 0.63 0.22 0.36 0.36 0.24 0.37
paraphrase_1 1.45 1.83 1.53 1.39 1.65 1.44 1.55
paraphrase_2 1.36 1.80 1.44 1.31 1.57 1.33 1.47
paraphrase_3 1.30 1.77 1.37 1.26 1.53 1.28 1.42
paraphrase_4 1.26 1.73 1.34 1.25 1.47 1.24 1.38
paraphrase_5 1.23 1.76 1.25 1.20 1.44 1.16 1.34
Table 14: Quality meta-evaluation of texts generated by different approaches. Mean scores for the two selected quality metrics are provided, averaged across the three generators (Aya, Falcon, and Opt).
The study of METAL (Hada et al., 2024) have evaluated usability of LLMs to be used as judges for evaluation of quality of the generated text in 10 languages. Although the primary focus of the study is on the summarization task, observation regarding generic text quality evaluation (i.e., the Linguistic Acceptability and Output Content Quality metrics) can be transferred to any text. The Linguistic Acceptability focuses more on a language structure alignment with the implicit norms and rules of a native speaker’s linguistic intuition. The Output Content Quality focuses more on relevance, clarity, originality, and linguistic fluency. The limitation of the study is in usage only of API-based "private" models, replicability of results of which is dependent on availability of the same versions via API and in ability of using seeds to make the output deterministic. There are also privacy and ethical concerns, when in some cases sensitive data just cannot be sent to API-based services due to policy restrictions. Beside scalability, one of the key benefits of meta-evaluation is its replicability (which is quite impossible in human evaluation) (Zhang et al., 2024). Therefore, we decided to use METAL dataset to evaluate correlation of various open LLMs to human judgements. The results are provided in Table 13, indicating that the SOTA open LLMs can be used for multilingual evaluation of text quality (comparable with GPT-4 performance for most languages).
Based on the results, we have selected Gemma-2, Qwen2, and Llama-3.1 as meta-evaluators for judging quality of the texts generated by the examined approaches. In total, 4012 text samples have been successfully evaluated by all three meta-evaluators. Inter-annotator agreement of the selected meta-evaluators is calculated in a form of pairwise Pearson correlation coefficient, averaging to 0.820.820.82. The definition of the selected Linguistic Acceptability and Output Content Quality metrics along with scoring schema (values of both being 0, 1, or 2, from lower to higher quality, respectively) can be found in METAL GitHub repository [^17] https://github.com/microsoft/METAL-Towards-Multilingual-Meta-Evaluation/tree/main/metrics/detailed. The meta-evaluation results of different approaches are summarized in Table 14, both metrics indicating that paraphrasing resulted in higher quality texts than the other two approaches, while each iteration of paraphrasing slightly reduces the text quality.
Test Language [mean]
Generator ar bg ca cs de el en es et ga gd hr hu nl pl pt ro ru sk sl uk zh $\rightarrow$ Average
Linguistic Acceptability
Aya-101 1.50 2.00 1.80 1.90 1.70 1.60 1.40 1.70 1.80 1.50 2.00 1.90 2.00 1.70 1.50 1.60 1.80 1.60 1.60 1.90 1.70 1.90 1.73
Gemini 2.00 1.80 1.70 1.90 1.70 1.90 2.00 1.80 2.00 1.40 1.50 2.00 1.50 1.70 1.80 1.90 2.00 2.00 1.90 1.90 2.00 1.90 1.83
GPT-3.5-Turbo-0125 1.50 1.90 1.60 1.80 1.40 1.50 1.70 1.50 1.20 1.70 1.30 1.80 1.50 1.40 1.20 1.50 1.40 2.00 1.80 1.60 1.60 2.00 1.59
Mistral-7B-Instruct-v0.2 1.20 1.90 1.70 1.60 1.10 1.10 2.00 1.80 0.80 1.80 1.90 1.50 1.60 1.60 1.40 1.90 1.20 1.60 1.50 1.70 1.70 1.50 1.55
OPT-IML-Max-30b 0.80 0.70 1.30 1.10 1.60 0.80 1.90 1.30 1.30 1.60 1.80 1.70 1.10 1.10 1.30 1.40 1.80 0.50 0.80 1.60 0.70 0.80 1.23
v5-Eagle-7B-HF 1.60 1.70 1.70 1.70 1.70 1.90 1.80 1.60 1.50 1.60 1.80 1.60 1.60 1.80 1.50 1.90 1.70 1.80 1.80 1.80 1.50 1.60 1.69
Vicuna-13b 1.30 1.50 1.90 1.30 1.80 1.00 2.00 2.00 0.80 1.20 1.80 1.40 1.40 1.70 1.60 1.90 1.80 1.60 1.20 1.50 1.70 1.70 1.55
human 1.60 1.00 1.00 0.60 1.40 0.70 1.00 0.90 0.60 1.70 1.60 0.80 0.50 0.80 1.30 1.00 1.20 1.10 0.80 1.70 1.20 1.80 1.10
Average 1.44 1.56 1.59 1.49 1.55 1.31 1.73 1.58 1.25 1.56 1.71 1.59 1.40 1.48 1.45 1.64 1.61 1.52 1.42 1.71 1.51 1.65 1.53
Output Content Quality
Aya-101 0.90 1.80 1.20 1.10 0.80 1.20 0.80 1.00 1.20 1.20 1.90 1.00 1.10 1.20 0.90 0.80 1.00 1.20 1.10 1.30 1.00 1.30 1.14
Gemini 1.90 1.70 1.40 1.60 1.30 1.90 1.90 1.80 1.80 1.20 1.50 2.00 1.40 1.70 1.50 1.80 1.70 1.90 1.90 1.80 1.90 1.70 1.70
GPT-3.5-Turbo-0125 1.20 1.50 1.20 1.20 1.10 1.00 1.10 1.20 0.70 1.20 1.30 1.10 1.10 1.30 0.80 1.10 1.10 1.30 1.40 1.40 1.00 1.70 1.18
Mistral-7B-Instruct-v0.2 0.60 1.30 1.30 1.10 1.10 0.80 1.70 1.60 0.70 1.20 1.40 1.20 1.20 1.20 1.10 1.50 0.70 1.10 1.10 1.40 1.20 1.30 1.17
OPT-IML-Max-30b 0.40 0.30 0.90 0.80 1.30 0.30 1.00 1.00 0.70 0.70 1.10 0.70 0.30 0.90 0.50 0.60 0.80 0.20 0.80 1.00 0.40 0.30 0.68
v5-Eagle-7B-HF 1.00 1.20 1.20 1.40 1.40 1.50 1.40 0.90 1.20 0.80 1.40 1.10 1.20 1.40 0.90 1.70 1.30 1.20 1.40 1.40 1.20 1.40 1.25
Vicuna-13b 0.70 1.00 1.30 0.90 1.50 0.70 1.60 1.40 0.50 0.90 0.90 0.80 1.10 1.30 1.10 1.20 1.00 1.20 1.00 1.30 1.20 1.20 1.08
human 1.20 0.30 0.60 0.30 1.10 0.40 0.60 0.60 0.30 1.20 0.80 0.50 0.30 0.50 0.30 0.70 0.70 0.80 0.60 1.30 0.80 1.40 0.70
Average 0.99 1.14 1.14 1.05 1.20 0.97 1.26 1.19 0.89 1.05 1.29 1.05 0.96 1.19 0.89 1.17 1.04 1.11 1.16 1.36 1.09 1.29 1.11
Table 15: Per-language quality meta-evaluation of texts generated by each generator. Mean of majority-voted (out of three meta-evaluators) scores of 10 samples for each combination are provided.
Similarly, we have used such meta-evaluation for quality assessment of the final texts generated by the selected generators. For this purpose, we have used a balanced subset of texts (10 samples per 22 languages per 7 generators and 1 human source, i.e. 1760 samples), resulting in 1752 evaluated samples (due to only 2 samples remained available from Gemini for Scottish Gaelic). The meta-evaluation scores given by the three meta-evaluators (pairwise Pearson correlation coefficient averaging to 0.690.690.69) are combined using the majority voting. The results, summarized in Table 15, indicate that the LLM generators generated texts of similar or higher quality across languages than the quality of original human texts. The reason might be an informal style used at social media (mistakes, spell errors, slang), which is more difficult for language models to follow (usually pre-trained on more formal web content). On average, the worst quality texts are generated by OPT-IML-Max-30B (failing mostly for non-Latin languages, still being on par with human text quality on average), while the best quality is provided by Gemini and Aya-101 models. Although not balanced across platforms (due to not each language being represented), meta-evaluation revealed the lowest quality of texts from Discord, followed by Telegram, and the highest quality of texts from Twitter.
B.5 Limited Bias Analysis
To minimize bias in the proposed dataset, we have run multiple existing detectors for data analysis. Based on the multilingual toxicity detector [^18] https://huggingface.co/textdetox/xlmr-large-toxicity-classifier (Dementieva et al., 2024), about 8% of the text samples are probably toxic (ranging from 5% in WhatsApp to 10% in Twitter parts). Based on the social media text topic detector [^19] https://huggingface.co/cardiffnlp/tweet-topic-latest-multi (Antypas et al., 2022), which is English-only, the topic distribution is illustrated in Figure 2. Based on the multilingual text genre detector [^20] https://huggingface.co/classla/xlm-roberta-base-multilingual-text-genre-classifier (Kuzman et al., 2023), the genre distribution is illustrated in Figure 3. Although the used detection cannot be considered thorough (using existing detectors in zero-shot manner cannot be considered fully accurate), when used just as an indication, we can see that the proposed dataset texts are distributed among various topics and genres; thus, limiting the presence of such a bias.
Figure 2: Detected topics in the MultiSocial dataset.
Figure 3: Detected genres in the MultiSocial dataset.
@@ Appendix D Cross-domain Evaluation
In Table 16 and 17, the results are provided in the same way as in Table 4, while testing on MULTITuDE (news domain) data. In pure cross-domain evaluation (Table 16), the machine texts are generated by the same generators (as used for training), in out-of-distribution evaluation (Table 17), the machine texts are generated only by Llama-2-70b (not available in MultiSocial for training).
@@ Appendix F Results Data
Tables 20-22 contain per-generator AUC ROC performance of each MGT detection method for each test language, Tables 23-25 contain per-platform AUC ROC performance of each MGT detection method for each test language, and Tables 26-28 contains per-generator AUC ROC performance of each MGT detection method for each platform, separately for each MGTD category. Table 29 contains cross-lingual evaluation of differently fine-tuned MGT detectors. Table 30 contains cross-platform evaluation of differently fine-tuned MGT detectors.
@@
BLOOMZ-3b-mixed-Detector
Aya-101 0.83 0.88 0.82 0.88 0.85 0.82 0.85 0.84 0.91 0.79 0.68 0.86 0.92 0.84 0.84 0.85 0.78 0.78 0.83 0.76 0.81 0.86 0.83
@@
ChatGPT-Detector-RoBERTa-Chinese
Aya-101 0.75 0.79 0.73 0.63 0.75 0.70 0.82 0.75 0.78 0.61 0.52 0.61 0.72 0.67 0.61 0.61 0.67 0.73 0.60 0.60 0.71 0.76 0.67
@@
RoBERTa-large-OpenAI-Detector
Aya-101 0.76 0.46 0.46 0.19 0.40 0.72 0.73 0.43 0.32 0.37 0.46 0.31 0.20 0.39 0.31 0.35 0.36 0.57 0.35 0.45 0.43 0.69 0.43
@@
ruRoBERTa-ruatd-binary
Aya-101 0.50 0.61 0.57 0.47 0.48 0.45 0.60 0.54 0.48 0.47 0.49 0.53 0.56 0.57 0.48 0.50 0.43 0.64 0.52 0.57 0.50 0.45 0.53
@@
XLM-RoBERTa-large-MultiSocial
Aya-101 0.89 0.97 0.94 0.97 0.95 0.94 0.93 0.94 0.95 0.81 0.73 0.94 0.98 0.95 0.95 0.91 0.93 0.90 0.96 0.87 0.86 0.85 0.93
@@
mDeBERTa-v3-base-MultiSocial
Aya-101 0.88 0.97 0.92 0.97 0.95 0.93 0.94 0.94 0.97 0.86 0.74 0.96 0.99 0.95 0.95 0.93 0.93 0.89 0.95 0.89 0.87 0.88 0.93
@@
Fast-Detect-GPT
Discord N/A N/A 0.64 0.81 0.81 N/A 0.86 0.82 0.70 0.76 0.76 0.81 0.79 0.81 0.83 0.84 0.80 N/A 0.65 N/A N/A N/A 0.79
@@
BLOOMZ-3b-mixed-Detector
Discord N/A N/A 0.96 0.89 0.87 N/A 0.87 0.90 0.87 0.80 0.66 0.84 0.86 0.91 0.91 0.89 0.78 N/A 0.71 N/A N/A N/A 0.87
@@
ChatGPT-Detector-RoBERTa-Chinese
Discord N/A N/A 0.88 0.72 0.86 N/A 0.95 0.88 0.85 0.73 0.62 0.66 0.82 0.71 0.70 0.75 0.83 N/A 0.68 N/A N/A N/A 0.77
@@
Longformer Detector
Discord N/A N/A 0.25 0.41 0.47 N/A 0.68 0.44 0.40 0.50 0.46 0.50 0.40 0.42 0.44 0.36 0.33 N/A 0.36 N/A N/A N/A 0.44
@@
RoBERTa-large-OpenAI-Detector
Discord N/A N/A 0.25 0.05 0.12 N/A 0.44 0.10 0.18 0.32 0.52 0.12 0.09 0.09 0.07 0.09 0.09 N/A 0.22 N/A N/A N/A 0.16
@@
ruRoBERTa-ruatd-binary
Discord N/A N/A 0.51 0.41 0.25 N/A 0.53 0.42 0.35 0.41 0.36 0.45 0.43 0.43 0.40 0.35 0.25 N/A 0.32 N/A N/A N/A 0.40
@@
BLOOMZ-3b-MultiSocial
Discord N/A N/A 0.99 1.00 0.97 N/A 0.99 0.99 0.99 0.97 0.84 0.98 1.00 0.98 0.99 0.99 0.99 N/A 0.97 N/A N/A N/A 0.99
@@
Falcon-rw-1b-MultiSocial
Discord N/A N/A 0.99 1.00 0.97 N/A 0.99 0.98 0.99 0.96 0.87 0.98 1.00 0.99 0.99 0.99 0.99 N/A 0.97 N/A N/A N/A 0.99
@@
Llama-3-8b-MultiSocial
Discord N/A N/A 0.99 1.00 0.97 N/A 0.99 0.99 1.00 0.96 0.88 0.99 1.00 0.99 1.00 0.99 0.99 N/A 0.98 N/A N/A N/A 0.99
@@
Mistral-7b-v0.1-MultiSocial
Discord N/A N/A 0.99 1.00 0.97 N/A 0.99 0.99 1.00 0.96 0.91 0.99 1.00 0.99 1.00 0.99 0.99 N/A 0.98 N/A N/A N/A 0.99
@@
XLM-RoBERTa-large-MultiSocial
Discord N/A N/A 0.96 0.99 0.96 N/A 0.98 0.97 0.99 0.89 0.72 0.98 1.00 0.98 0.99 0.97 0.98 N/A 0.99 N/A N/A N/A 0.97
@@
mDeBERTa-v3-base-MultiSocial
Discord N/A N/A 0.98 1.00 0.97 N/A 0.98 0.98 0.99 0.91 0.75 0.98 1.00 0.98 0.99 0.99 0.99 N/A 0.97 N/A N/A N/A 0.98
@@
Binoculars
Mistral-7B-Instruct-v0.2 0.77 0.62 0.68 0.68 0.67 0.68
@@
DetectLLM-LRR
Mistral-7B-Instruct-v0.2 0.94 0.63 0.71 0.71 0.68 0.71
@@
Fast-Detect-GPT
Mistral-7B-Instruct-v0.2 0.65 0.51 0.58 0.59 0.62 0.58
@@
LLM-Deviation
Mistral-7B-Instruct-v0.2 0.95 0.63 0.69 0.70 0.67 0.70
@@
BLOOMZ-3b-mixed-Detector
Mistral-7B-Instruct-v0.2 0.87 0.62 0.76 0.70 0.78 0.74
@@
ChatGPT-Detector-RoBERTa-Chinese
Mistral-7B-Instruct-v0.2 0.85 0.70 0.75 0.74 0.85 0.76
@@
Longformer Detector
Mistral-7B-Instruct-v0.2 0.51 0.55 0.59 0.44 0.47 0.52
@@
RoBERTa-large-OpenAI-Detector
Mistral-7B-Instruct-v0.2 0.11 0.38 0.30 0.38 0.33 0.31
@@
ruRoBERTa-ruatd-binary
Mistral-7B-Instruct-v0.2 0.47 0.52 0.54 0.53 0.53 0.53
@@
Aya-101-MultiSocial
Mistral-7B-Instruct-v0.2 1.00 0.98 0.99 0.99 1.00 0.99
@@
BLOOMZ-3b-MultiSocial
Mistral-7B-Instruct-v0.2 1.00 0.94 0.97 0.98 0.99 0.97
@@
Falcon-rw-1b-MultiSocial
Mistral-7B-Instruct-v0.2 1.00 0.93 0.97 0.98 0.98 0.97
@@
Llama-3-8b-MultiSocial
Mistral-7B-Instruct-v0.2 1.00 0.98 0.99 0.99 1.00 0.99
@@
Mistral-7b-v0.1-MultiSocial
Mistral-7B-Instruct-v0.2 1.00 0.97 0.99 0.99 1.00 0.99
@@
XLM-RoBERTa-large-MultiSocial
Mistral-7B-Instruct-v0.2 0.99 0.95 0.97 0.97 0.99 0.97
@@
mDeBERTa-v3-base-MultiSocial
Mistral-7B-Instruct-v0.2 0.99 0.94 0.98 0.98 0.99 0.97
@@
XLM-RoBERTa-large
Discord 0.99 0.87 0.90 0.83 0.94 0.90
@@
mDeBERTa-v3-base
Discord 0.99 0.86 0.89 0.84 0.93 0.89